You have $2000 that you invested in an account t today. Interest is compounded monthly and you will withdaw $4000 in exactly 10 years. What is the nominal interest rate?
Answer #1
We have the equality,
4000 = 2000*(1+r)^120, where r = nominal interest rate/12.
Solving for r,
r = (4000/2000)^(1/120)-1 = 0.005792941
Nominal interest rate = r*12 = ((4000/2000)^(1/120)-1)*12 = 6.95%
CHECK:
FV = 2000*(1+0.0695/12)^120 = $        3,999.39
Difference due to approximation in r*12